# Portable, Massively Parallel Implementation of a Material Point Method for Compressible Flows

### v2

@@ Abstract
The recent evolution of software and hardware technologies is leading to a renewed computational interest in Particle-In-Cell (PIC) methods such as the Material Point Method (MPM). Indeed, provided some critical aspects are properly handled, PIC methods can be cast in formulations suitable to the requirements of data locality and fine-grained parallelism of modern hardware accelerators as Graphics Processing Units (GPUs). Such a rapid and continuous technological development increases also the importance of generic and portable implementations. While continuum mechanics simulations have already shown the capabilities of MPM on a wide range of phenomena, the use of the method in compressible fluid dynamics is less frequent. In this paper we present a portable, highly parallel, GPU based MPM solver for compressible gas dynamics. The implementation aims to reach a good compromise between portability and efficiency and to give a first assessment of the potential of this approach in reproducing strongly compressible gas flows, also taking into account solid obstacles. The proposed model constitutes a first step towards the development of a monolithic MPM solver for Fluid-Structure Interaction (FSI) problems at all Mach numbers up to the supersonic regime.
@@
This work has not yet been peer-reviewed and is provided by the contributing authors as a means to ensure timely dissemination of scholarly and technical work on a noncommercial basis. Copyright and all rights therein are maintained by the authors or by other copyright owners. It is understood that all persons copying this information will adhere to the terms and constraints invoked by each author’s copyright. This work may not be reposted without explicit permission of the copyright owner. This work has been submitted for publication. Copyright in this work may be transferred without further notice.
@@ 1 Introduction and Motivation
Hardware and software technology evolution has led to a renewed interest in particle-based numerical techniques for differential problems, such as the Material Point Method (MPM, [51, 63]), that was developed stemming from the Particle In Cell method Fluid Implicit Particle Method [29, 8] to treat history-dependent materials, thanks to their suitability to modern hardware accelerators requirements [28, 22]. Indeed, parallel computing is a key capability for effective advancement of industrial-grade, large-scale computations, particularly in the fields of high-speed, compressible computational fluid dynamics and Fluid-Structure Interaction (FSI, [37]) simulations. On the other hand, rapid hardware and software evolution makes researchers lean towards generic and portable implementations, which enables developers to adapt fast to sudden technology innovations.
In view of exponentially increasing High Performance Computing (HPC) resources and the method’s accurate handling of a wide range of phenomena, MPM is becoming increasingly relevant for compressible CFD and FSI simulations. In MPM, continuum media are discretised by means of Lagrangian particles, the material points, each one storing every state variable evaluated at its position, enabling the description of a wide range of materials [26, 49, 32] as well as large deformations [1, 58, 18] and fracture phenomena [60]. Notwithstanding its Lagrangian character, MPM also uses a background Cartesian grid to compute differential quantities and solve the motion equation, thus mediating the particle-particle interactions, and taking advantage of both Eulerian and Lagrangian approaches. These features make it an interesting candidate for the development of a highly parallel, Graphics Processing Units (GPUs)-based FSI solver. Indeed, in order to reach a high degree of efficiency, GPU-based solvers require a large degree of data locality and fine-level parallelism. MPM implementations can naturally achieve these, since:
@@
Since MPM first introduction, various enhancements of the method were proposed, solving some of its drawbacks such as grid-crossing error [55] and volumetric locking [17], and improving its accuracy and conservation properties. To address instabilities which arise when particles cross the grid, variants of the method were proposed, such as GIMP [4], CPDI [45], iMPM [52] and IGA-MPM [40], and numerical analysis studies on the use of B-Splines as basis functions in MPM were conducted [48]. An affine and a polynomial mapping was later added to the B-Splines MPM [33, 23] resulting in an angular momentum-conserving model. Other studies explored least squares techniques [62] enabling the recovery of the affine/polynomial splines MPM [31], the effect of spatial and temporal discretisation errors, and the application of symplectic integrators [5]. In light of these developments, MPM represents an effective method for the simulation of a wide range of materials and phenomena.
However, in spite of recent implementations for computer graphics applications [24, 61], MPM implementations on GPUs have not yet been applied to compressible fluid dynamics. In fact, even when addressing gas dynamics and solids mechanics, MPM has been used only for the solids, while the compressible fluid dynamic problem has been solved through a second order WENO scheme [12]; moreover, such implementation is run on CPUs. Existing studies on MPM for compressible flows focused on numerical convergence and error analysis [57] and on modelling and expanding the range of applications for the method [65, 50]. Nonetheless, to the best of the authors’ knowledge, no studies exist that are dedicated to the GPU based, High Performance Computing aspects of MPM implementations for strongly compressible fluid dynamics problems.
@@
Therefore, existing MPM implementations either present challenges in adapting to modern hardware or are not widely studied and employed in compressible fluid dynamics. A good example is given by high-order WENO and ENO schemes, that guarantee numerically excellent results in gas dynamics. For these schemes to work properly, the stencil should be able to grow in any direction into neighbouring mesh elements. From a computational point of view, this works very well on structured grids. However, designing proper structured grids for FSI is significantly time consuming. On the other hand, if unstructured grids are chosen, growing stencils lead to a large increase of halo transfers in communication among parallel processors, thus requiring a large consumption of memory bandwidth and generating a large amount of indirection, degrading the suitability of the methods for modern parallel hardware. Moreover, such schemes have to be coupled with appropriate methods for structural mechanics, that have to be suitable for modern parallel hardware too.
@@
  • it benefits from the choice of employing structured Cartesian grids, since continua discretisation is delegated to Lagrangian material particles, allowing the modelling of large deformations as well;
@@
However, even if MPM stemmed from fluid solvers, there are only a few studies addressing compressible fluid dynamics with MPM. In addition, the existing studies, some of which (e.g. [65]) predate work on numerical linear algebra on GPUs (e.g. [6, 35]), do not deal with parallelisation on modern hardware. In addition, the MPM community has preferentially paid attention to continuum mechanics modelling.
With this work we aim to address only the fluid portion of a future FSI solver, i.e., the portion which is usually left to Eulerian CFD solvers. We focus on algorithm and data structures suitable for implementation on various parallel architectures, with a particular attention to HPC GPUs, and we address the problem of reducing limitations in treating obstacles that derives from the choice of a structured Cartesian grid, by means of a signed distance function approach.
The current study focuses on linear MPM to better study its scalability features, leaving the update to a higher order MPM for future work. However, we anticipate a discussion on how higher regularity base functions could fit our approach and how the parallel efficiency would be influenced by the extension of base function support on more elements. From the modelling point of view, to give a first assessment of MPM in simulating efficiently compressible gases on GPUs, we treat the gas as inviscid fluid, neglecting the effects of turbulence, and adopt an artificial viscosity based shock capturing technique.
In previous work [2] we ported an existing MPM algorithm [65, 64, 14] to CUDA-C [41], adapting it to GPUs architecture and obtaining promising speed-ups with respect to a C++ CPU version. The CUDA-C programming model was chosen because of its known performance capabilities on NVIDIA A100 GPUs [42, 9]. However, the model had shortcomings in its low-level character and specificity to the underlying architecture. In addition, while it can be made to work on, e.g., AMD GPUs via HIP with minimal changes, it cannot be ported to different architectures such as CPUs without substantial code modifications.
One of NVIDIA’s solutions for performance portability involves its own C++ compiler nvc++ [21, 43], using which one rewrites algorithm steps relying on C++ Standard Template Library (STL), specifying a parallel execution policy. To give an example, in this framework it is possible to parallelise the following for loop:
@@
Although the approach does not prevent the developer from introducing data-races or deadlocks, the code can be written only once, and then, depending on the option specified during compilation, run in parallel on multi-core CPUs or GPUs.
However, nvc++ high-level features come at a price, as the developer loses the control on memory handling and, in particular, on when Host-to-Device and Device-to-Host memory copies are performed. The code runs on host, any invocation to an algorithm with a parallel execution policy expresses the developer’s preference to run that algorithm in parallel. If compiled for GPUs, memory is copied to the device, computations are performed and then data is copied back to host. In physics and in engineering dynamics simulations, where each time step depends on the previous one, this means that the maximum reachable level of parallelisation corresponds to a single call to a parallel algorithm inside a traditional for loop for time advancing. Since memory fetches can be significantly more costly than computations, this is sub-optimal.
The present contribution expands the scope of [2] and overcomes the performance issues with nvc++ by designing a portable MPM code for transonic and supersonic compressible gas dynamics based on Thrust [44], a C++ template library for CUDA based on the Standard Template Library (STL). In addition to nvc++ features, Thrust introduces two types of STL like vectors, one for the host (typically, the CPU) and one for the device (typically, the GPU); this allows the user to specify when memory should be copied via explicit calls. For example, it’s possible to perform a single Host to Device copy when reading the initial data file, and one Device To Host copy whenever saving partial results on output data files, strongly reducing execution time related to memory fetches.
@@
Notwithstanding the portability objectives, we designed the code for optimisation on modern graphics hardware accelerators and made further modifications to the original algorithm [14, 65]. For example, we:
  • added treatment of obstacles in gas dynamics simulations,
  • added a particle re-ordering step,
  • postponed particles movement to the end of the time loop,
and thus values of basis function evaluated at previous particle positions don’t need to be stored and can be computed on the fly instead, which is cheaper on modern hardware accelerators.
Specific development choices are described that enable optimisation on modern GPU-based HPC architectures and adaptability to different hardware, as well as generality of the approach for the purposes of its portability also towards different software. The accuracy and performance of the implementation are evaluated on a number of compressible flow benchmarks and results are compared with available approaches in the literature.
The paper is structured as follows. Section 2 contains the analytical and numerical formulation of the method. Section 3 describes the chosen algorithm and the current and implementation choices. Section 4 illustrates the results obtained on benchmark test cases, and the final section 5 contains a discussion of the results, draws the conclusions of the performed work, and prospects future developments.
@@ 2.1 Model Equations and Weak Formulation
As mentioned, MPM takes advantages of both the Eulerian and the Lagrangian perspective; specifically, the Lagrangian particles discretisation approach affords flexibility in modelling various kind of phenomena and materials, while the Eulerian grid discretisation is beneficial for scalability as the grid allows the mediation of particle-particle interactions. For the mathematical modelling we choose the Lagrangian framework, and we model compressible fluids in a domain $\Omega$ governed by the conservation laws for mass, linear momentum and energy densities [27, 64, 19]:
$\begin{cases}\dfrac{d\rho}{dt}+\rho\nabla\cdot\mathbf{v}=0\\[5.69054pt] \rho\dfrac{d\mathbf{v}}{dt}=\nabla\cdot\bm{\sigma}+\rho\mathbf{b}\\[5.69054pt] \rho\dfrac{de}{dt}=\bm{\sigma}:\bm{\dot{\varepsilon}}-\nabla\cdot\mathbf{q}+\rho s\end{cases}$ (1)
where $\rho$, $\mathbf{v}$, $\bm{\sigma}$, $\rho\mathbf{b}$, $p$, $e$, $\bm{\dot{\varepsilon}}$, $\mathbf{q}$ and $s$ represent respectively the mass density, the velocity, the stress tensor, the external forces density, the pressure, the specific internal energy, the strain rate tensor of the fluid, the heat flux and the heat source.
In this work, we:
  • enforce mass conservation by keeping constant the number of Lagrangian material points (particles) along the simulations, as well as each particle mass,
  • choose to model the fluid as a bi-atomic perfect gas, governed by the state equation
@@
    where $e$ is the specific internal energy and $\gamma$ is the ratio between specific heat at constant pressure and constant volume, $\gamma=c_{p}/c_{v}=1.4$ for a bi-atomic gas.
To obtain the final discretisation of the equations, a weak form of the force balance expressed by the motion equation is derived [51, 50]. Having chosen a proper functional space $V$ of regular enough functions, we multiply the momentum conservation equation in (1) by an arbitrary $\mathbf{w}\in V$, integrate over the fluid domain $\Omega$ and integrate by parts the divergence term.
The weak formulation of the problem then reads:
Find $\mathbf{v}$ such that, $\forall\mathbf{w}\in V$
@@
where $\mathbf{\dot{v}}$ denotes the total, or Lagrangian, velocity derivative, $\mathbf{w}$ is a generic test function and $\bm{\tau}$ is the stress on the surface. In the following, we simplify the model assuming the fluid to be inviscid, and we neglect external forces, heat sources and heat flux. We thus get the Euler equations for an inviscid, compressible fluid:
$\begin{cases}\dfrac{d\rho}{dt}+\rho\nabla\cdot\mathbf{v}=0\\[5.69054pt] \rho\dfrac{d\mathbf{v}}{dt}=-\nabla p\\[5.69054pt] \rho\dfrac{de}{dt}=-p\nabla\cdot\mathbf{v}\end{cases}$ (4)
Note that, in the above expression, the fluid stress tensor is simplified from:
$\sigma_{ij}=2\mu\dot{\varepsilon}_{ij}-\tfrac{2}{3}\mu\varepsilon_{kk}\delta_{ij}-p\delta_{ij}$
to:
$\sigma_{ij}=-p\delta_{ij}$
which greatly simplifies the energy conservation equation as well as the momentum equation.
Finally, following [63], we make use of a combination of von Neumann-Richtmyer [59] and Landshoff [36] artificial viscosity as shock-capturing scheme, relying on numerical damping to handle numerical oscillations in the vicinity of discontinuities, and computing them as part of the solution to problem (4). Specifically, application of this artificial viscosity-based shock-capturing method involves the addition of an artificial term $q$ to pressure $p$, substituting the pressure $p$ to the sum $p+q$ in problem (4). The artificial term is added only in compression regions, identified by the local velocity divergence being less than zero. Its expression reads
$q=\begin{cases}c_{0}\left(\rho\,h\,\mathrm{div}(v)\right)^{2}-c_{1}\,h\,\sqrt{\gamma\tfrac{p}{\rho}}\,\mathrm{div}(v)\,\,&\mathrm{div}(v)<0\\ 0\quad&\mathrm{div}(v)\geq 0\end{cases}$ (5)
where $c_{0},c_{1}$ are dimensionless constants and $h$ is the characteristic length of a grid cell. The effects of the two components of $q$ are shown in section 4.
This dynamically comprehensive yet relatively simple equation set is chosen as it enables testing the performance portability of the MPM implementation on compressible gas dynamics benchmarks [56, 47] and comparison with existing approaches [65, 50].
The continuity and energy conservation equations will be directly discretised and used to update particle properties in the algorithm (see section 3 and algorithm 1), while the discretisation of the weak form of the motion equation is related to the link between the Eulerian grid and the Lagrangian particles, and is treated in more depth in the next section.

### v4

@@ Abstract
The recent evolution of software and hardware technologies is leading to a renewed computational interest in Particle-In-Cell (PIC) methods such as the Material Point Method (MPM). Indeed, provided some critical aspects are properly handled, PIC methods can be cast in formulations suitable for the requirements of data locality and fine-grained parallelism of modern hardware accelerators such as Graphics Processing Units (GPUs). Such a rapid and continuous technological development increases also the importance of generic and portable implementations. While the capabilities of MPM on a wide range continuum mechanics problem have been already well assessed, the use of the method in compressible fluid dynamics has received less attention. In this paper we present a portable, highly parallel, GPU based MPM solver for compressible gas dynamics. The implementation aims to reach a good compromise between portability and efficiency in order to provide a first assessment of the potential of this approach in solving strongly compressible gas flow problems, also taking into account solid obstacles. The numerical model considered constitutes a first step towards the development of a monolithic MPM solver for Fluid-Structure Interaction (FSI) problems at all Mach numbers up to the supersonic regime.
@@
This preprint is a preliminary version of a paper published in Computational Particle Mechanics, doi: https://doi.org/10.1007/s40571-024-00864-2. URL of the published version: https://link.springer.com/article/10.1007/s40571-024-00864-2.
@@ 1 Introduction and Motivation
Hardware and software technology evolution has led to a renewed interest in particle-based numerical techniques for differential problems, such as the Material Point Method (MPM, [57, 68]), that was developed stemming from the Particle In Cell method Fluid Implicit Particle Method [29, 7] to treat history-dependent materials, thanks to their suitability to modern hardware accelerators requirements [28, 22]. Indeed, parallel computing is a key capability for effective advancement of industrial-grade, large-scale computations, particularly in the fields of high-speed, compressible computational fluid dynamics and Fluid-Structure Interaction (FSI, [39]) simulations. On the other hand, rapid hardware and software evolution makes researchers lean towards generic and portable implementations, which enable developers to swiftly adapt to sudden technology innovations.
In view of exponentially increasing High Performance Computing (HPC) resources and the method’s accurate handling of a wide range of phenomena, MPM is becoming increasingly relevant for compressible CFD and FSI simulations. In MPM, continuum media are discretised by means of Lagrangian particles, the material points, each one storing every state variable evaluated at its position, enabling the description of a wide range of materials [26, 55, 32] as well as large deformations [1, 63, 18] and fracture phenomena [65]. Notwithstanding its Lagrangian character, MPM also uses a background Eulerian grid, that is often chosen to be a tensor-product Cartesian grid for sake of efficiency, to compute differential quantities and solve the motion equation, thus mediating the particle-particle interactions, and taking advantage of both Eulerian and Lagrangian approaches. These features make MPM an interesting candidate for the development of a highly parallel, Graphics Processing Units (GPUs)-based FSI solver. Indeed, in order to reach a high degree of efficiency, GPU-based solvers require a large degree of data locality and fine-level parallelism, and MPM implementations can naturally achieve these, since:
@@
Since the first introduction of MPM, various enhancements of the method were proposed, solving some of its drawbacks such as grid-crossing instabilities [60] and volumetric locking [16], and improving its accuracy and conservation properties. To address instabilities which arise when particles cross the grid, variants of the method were proposed, such as GIMP [4], CPDI [48], iMPM [58] and IGA-MPM [43], and numerical analysis studies on the use of B-Splines as basis functions in MPM were conducted [53]. An affine and a polynomial mapping were later added to the B-Splines MPM [33, 23] resulting in an angular momentum-preserving models. Other studies explored least squares techniques [67] enabling the recovery of the affine/polynomial splines MPM [31], the effect of spatial and temporal discretisation errors, and the application of symplectic integrators [5]. In light of these developments, MPM represents an effective method for the simulation of a wide range of materials and phenomena.
However, even though MPM was actually initially developed for fluid simulations, there are only a few recent studies addressing compressible fluid dynamics using MPM.
Existing studies on MPM for compressible flows focused on modelling, as the first work addressing compressible fluids and thin membranes [70], on numerical convergence and error analysis [62], and on expanding the range of applications for the method, as the more recent paper [56], which also takes into account thermomechanics effects; finally, in [34] the use of PolyPIC transfers [23] to construct a second order accurate discretisation of the Navier-Stokes equations within a PIC-MPM context is proposed. Nonetheless, none of these studies addresses high speed gas dynamics. In addition, the existing studies, some of which (e.g. [70]) predate work on numerical linear algebra on GPUs (e.g. [6, 37]), do not deal with parallelisation on modern hardware.
In fact, in recent years MPM has gained popularity in the computer graphics community (e.g. [54]), leading to relevant results on modern hardware. In particular, Gao et al. [24] designed an MPM implementation optimised for GPUs, adapting the efficient data structure for sparse data on CPU of Setaluri et al. ([50]), and Wang et al. [66] extended the work to a multi-GPU shared memory, MPI-free programming model. Finally, Buckland et al. [10] recently considered the problem of portability of a parallel MPM implementation in the computational mechanics realm.
However, these works do not treat compressible fluid dynamics problems, and the choice of the material to simulate has additional consequences on the computational efficiency.
For example, the GPU-optimised SPGrid data structure of [24] is designed for a computational domain with a relevant portion of empty cells, as it employs MPM to simulate solids, elasto-plastic materials, and sand. Conversely, gases fill the computational domain, requiring a different data structure. The efficiency of MPM on modern, parallel hardware in the modelling of compressible gas dynamics has yet to be proved. Indeed, existing studies addressing Fluid-Structure Interaction in a compressible, high speed gas dynamics context typically rely on a segregated approach, coupling different solvers for different materials. This approach, besides the known pros and cons with respect to monolithic solvers, also introduces challenges in the performance portability of the implementation.
@@
A very good example of this approach is given by [11], where MPM has been used only for the solids, while the compressible fluid dynamic problem has been solved through a second order WENO scheme, and the code is run on CPUs. Indeed, high-order WENO and ENO schemes guarantee numerically excellent results in gas dynamics but, for these schemes to work properly, the stencil should be able to grow in any direction into neighbouring mesh elements. From a computational point of view, this works very well on structured grids. However, designing proper structured grids for Eulerian methods in FSI is significantly time consuming. On the other hand, if unstructured grids are chosen, growing stencils leads to a large increase of halo transfers in communication among parallel processors, thus requiring a large consumption of memory bandwidth and generating a large amount of indirection, degrading the suitability of the methods for modern parallel hardware.
@@
  • it benefits from the choice of employing structured, Cartesian tensor product grids, since continua discretisation is delegated to Lagrangian material particles, allowing the modelling of large deformations without the need of complex and costly meshes;
@@
Nonetheless, to the best of the authors’ knowledge, no studies exist that are dedicated to the GPU based, High Performance Computing aspects of MPM implementations for strongly compressible fluid dynamics problems. Our aim is to explore these aspects in depth, considering also the performance portability features of the algorithm and its implementation.
With this work we aim to address only the fluid portion of a future FSI solver, i.e., the portion which is usually left to Eulerian CFD solvers. We focus on algorithm and data structures suitable for implementation on various parallel architectures, with a particular focus on HPC GPUs, and we address the problem of reducing limitations in treating obstacles that derives from the choice of a structured Cartesian grid, by means of a signed distance function approach.
The current study focuses on linear MPM to better study its scalability features, leaving the update to a higher order MPM for future work. However, we anticipate a discussion on how higher regularity base functions could fit our approach and how the parallel efficiency would be influenced by the extension of base function support on more elements. From the modelling point of view, to give a first assessment of MPM in simulating efficiently compressible gases on GPUs, we treat the gas as inviscid fluid, neglecting the effects of turbulence, and adopt an artificial viscosity based shock-capturing technique.
In previous work [2], we ported an existing MPM algorithm [70, 69, 13] to CUDA-C [44], adapting it to GPU architecture and obtaining promising speed-ups with respect to a C++ CPU version. The CUDA-C programming model was chosen because of its known performance capabilities on NVIDIA A100 GPUs [45, 8]. However, the programming model had shortcomings in its low-level character and specificity to the underlying architecture: while it can be made to work on, e.g., AMD GPUs via HIP with minimal changes, it cannot be ported to different architectures such as multicore CPUs without substantial code modifications.
One of NVIDIA’s solutions for performance portability involves its own C++ compiler nvc++ [21, 46], using which, one rewrites algorithm steps relying on C++ Standard Template Library (STL), specifying a parallel execution policy. To give an example, in this framework it is possible to parallelise the following for loop:
@@
Although the approach does not prevent the developer from introducing data-races or deadlocks, code can be written only once, and then, depending on options specified at compile time, run in parallel on multi-core CPUs or GPUs.
However, nvc++ high-level features come at a price, as the developer loses the control on memory handling and, in particular, on the moment when Host-to-Device and Device-to-Host memory copies are performed. The code runs on host, any invocation to an algorithm with a parallel execution policy expresses the developer’s preference to run that algorithm in parallel. If compiled for GPUs, memory is copied to the device, computations are performed and then data is copied back to host. In physics and in engineering dynamics simulations, where each time step depends on the previous one, this means that the maximum reachable level of parallelisation corresponds to a single call to a parallel algorithm inside a traditional for loop for time advancing. Since memory fetches can be significantly more costly than computations, this is sub-optimal.
The present contribution expands the scope of [2] and overcomes the performance issues with nvc++ by designing a portable MPM code for transonic and supersonic compressible gas dynamics based on Thrust [47], a template library for CUDA and HIP based on the C++ Standard Template Library (STL). In addition to nvc++ features, Thrust introduces two types of STL like vectors, one for the host (typically, the CPU) and one for the device (typically, the GPU); this allows the user to specify when memory should be copied via explicit calls. For example, it is possible to perform a single Host to Device copy when reading the initial data file, and one Device To Host copy whenever saving partial results on output data files, strongly reducing execution time related to memory fetches.
@@
  • Thrust itself allows for modifications to the default device and host system, so that it can be run both serially and in parallel, with TBB or OpenMP on CPUs for the host, and also on the GPU via CUDA-C++ for the device only. The code can be compiled and executed on a CPU-only machine as well, provided THRUST_DEVICE_SYSTEM is defined to be
@@
Specific development choices are described that enable optimisation on modern GPU-based HPC architectures and portability to different hardware, as well as generality of the approach for the purposes of its adaptability also towards different software. The accuracy and performance of the implementation are evaluated on a number of compressible flow benchmarks and the results are compared with available approaches in the literature.
The paper is structured as follows. Section 2 contains the analytical and numerical formulation of the method. Section 3 describes the chosen algorithm and the current and implementation choices. Section 4 illustrates the results obtained on benchmark test cases, addressing numerical modelling aspects as well as performance portability, and the final section 5 contains a discussion of the results, draws the conclusions of the performed work, and prospects future developments.
@@ 2 Model Formulation
In this section we present, in (5) of subsection 2.1, the model equations that are actually solved by our code. In order to clearly point out the simplifications that we adopted and the physical limitations they correspond to, we briefly outline the derivation of such model from the more general form stated in (1).
As previously mentioned, MPM takes advantages of both the Eulerian and the Lagrangian perspective; specifically, the Lagrangian particles discretisation approach affords flexibility in modelling various kind of phenomena and materials, while the Eulerian grid discretisation is beneficial for scalability as the grid allows the mediation of particle-particle interactions.
As the Eulerian approach used for the integration of the moment conservation equation is based on the weak form of such equation, the weak formulation is presented at the end of subsection 2.1, before proceeding to the description of the numerical algorithm in subsection 2.2
@@ 2.1 Model Equations and Weak Formulation
Denoting by $\dfrac{d\cdot}{dt}$ the material derivative operator, the flow of a compressible fluid characterised by its mass density $\rho$, velocity $\mathbf{v}$ and internal energy $e$ and contained in a domain $\Omega$ is governed by the following set of conservation laws  [27, 69, 19]:
$\begin{cases}\dfrac{d\rho}{dt}+\rho\nabla\cdot\mathbf{v}=0\\[5.69054pt] \rho\dfrac{d\mathbf{v}}{dt}=\nabla\cdot\bm{\sigma}+\rho\mathbf{b}\\[5.69054pt] \rho\dfrac{de}{dt}=\bm{\sigma}:\bm{\dot{\varepsilon}}-\nabla\cdot\mathbf{h}+\rho s.\end{cases}$ (1)
In the equations above, $\bm{\sigma}$, $\rho\mathbf{b}$, $\bm{\dot{\varepsilon}}$, $\mathbf{h}$ and $s$ represent the stress tensor, the external force density, the strain rate tensor, the heat flux and the heat source, respectively; suitable constitutive relations for such quantities must be provided in order for the model to be solvable, our choices for such relations are described below. For the stress tensor $\bm{\sigma}$ we assume the following Ansatz
$\sigma_{\alpha\beta}=-(p+q)\ \delta_{\alpha\beta}$ (2)
where $\delta_{\alpha\beta}$ is the Kronecker delta.
In (2), $p$ denotes the thermodynamic pressure which we assume to be related to the state variables via the state equation for a perfect gas
@@
where $\gamma$ is the ratio between specific heat at constant pressure and constant volume which, for a bi-atomic perfect gas, is given by $\gamma=c_{p}/c_{v}=\tfrac{7/2}{5/2}=1.4$. While viscous stresses are neglected in the expression of (2) (indeed we focus on inviscid fluid flow only in the present work), a small artificial viscosity is commonly adopted in numerical simulations to improve stability. In this paper this is accomplished, as suggested in [68], via the additional pressure term $q$ in (2) which consists of a combination of von Neumann-Richtmyer [64] and Landshoff [38] artificial viscosities and is meant to damp out spurious oscillations near solution discontinuities. The artificial viscosity correction is only activated in compression regions, identified as areas where the divergence of $\mathbf{v}$ is less than zero, and its expression reads
$q=\begin{cases}c_{0}\left(\rho\,h\,\mathrm{div}(\mathbf{v})\right)^{2}-c_{1}\,h\,\sqrt{\gamma\tfrac{p}{\rho}}\,\mathrm{div}(\mathbf{v})\,\,&\mathrm{div}(\mathbf{v})<0\\ 0\quad&\mathrm{div}(\mathbf{v})\geq 0\end{cases}$ (4)
where $c_{0},c_{1}$ are dimensionless constants and $h$ denotes the characteristic length of a grid cell. In section 4 we assess the effect of the different terms comprising the correction $q$ on the solution of a problem displaying a shock wave, a contact discontinuity and a rarefaction wave. Finally, we consider an adiabatic, non reactive flow regime and, therefore, neglect heat fluxes and sources, as well as external body forces.
In conclusion, the system of equations we focus on in the present study can be summarised as
$\begin{cases}\dfrac{d\rho}{dt}+\rho\nabla\cdot\mathbf{v}=0\\[5.69054pt] \rho\dfrac{d\mathbf{v}}{dt}=-\nabla(p+q)\\[5.69054pt] \rho\dfrac{de}{dt}=-(p+q)\nabla\cdot\mathbf{v}\\[5.69054pt] \end{cases}$ (5a)
and
$\small\begin{cases}p=(\gamma-1)\rho e\\[5.69054pt] q=c_{0}\left(\rho\,h\,\mathrm{div}(\mathbf{v})\right)^{2}-c_{1}\,h\,\sqrt{\gamma\tfrac{p}{\rho}}\,\mathrm{div}(\mathbf{v}),&\mathrm{div}(\mathbf{v})<0\\ q=0,&\mathrm{div}(\mathbf{v})\geq 0.\end{cases}$ (5b)
Of course (5) must be supplemented by suitable initial and boundary conditions, which will be detailed separately for each of the individual flow configurations considered in the following.
This dynamically comprehensive yet relatively simple equation set is chosen as it enables testing the performance portability of the MPM implementation on compressible gas dynamics benchmarks [61, 51] and comparison with existing approaches [70, 56] (see eqs. (2.55)-(2.61) in [68] for a comparison).
As shown in the next section, the momentum equation is solved on the grid, while mass conservation, internal energy conservation and state equations are solved on particles (see step 10 in 1.). The link between the Lagrangian particles and the Eulerian mesh is given by mappings which can be derived from the weak formulation of the motion equation [57, 68], that is obtained as follows.
Let $\Omega\in\mathbb{R}^{2}$ be a domain, and $\mathbf{\dot{v}}:\Omega\rightarrow\mathbb{R}^{2}$ denote the total, material, derivative of the velocity. Then, momentum conservation reads
$\rho\mathbf{\dot{v}}=-\nabla(p+q)$ (6)
Take a test function $\mathbf{w}\in V=[H_{0}^{1}(\Omega)]^{2}$, multiply (6) by $\mathbf{w}$ and integrate over the domain $\Omega$
@@
Integrating by parts the second term
[…]
where boundary integrals are neglected because of the requirement that $\mathbf{w}\in V=[H_{0}^{1}(\Omega)]^{2}$.
The continuity, the energy conservation and the state equations will be directly discretised and used to update particle properties in the algorithm (see section 3 and algorithm 1), while the discretisation of the weak form of the momentum equation, related to the link between the Eulerian grid and the Lagrangian particles, is treated in more depth in the next section.